# Breaking Boundaries: Investigating the Effects of Model Editing on Cross-linguistic Performance

### v3

@@ 1 Introduction
Handling multilinguality in language models remains a significant challenge, particularly when models are prompted in languages other than English. Tasks such as question answering Xu et al. (2024), addressing multilingual safety concerns Wang et al. (2024); Deng et al. (2024), or performing knowledge edits Hazra et al. (2024) often reveal noticeable gaps in performance for low-resource languages. Despite the advancements in multilingual large language models (LLMs), disparities persist, especially for languages with fewer computational resources. A clear example of this issue arises in knowledge editing Sinitsin et al. (2020); De Cao et al. (2021). For instance, when an LLM is updated to correct a factual statement, “The PM of the UK is Rishi Sunak" to “The PM of the UK is Keir Starmer" the model may apply the update accurately in well-represented languages like English or French Qi et al. (2023); Xu et al. (2023). However, the same edit often fails to propagate when queried in low-resourced languages like Tamil or Hindi. This inconsistency highlights a critical weakness in the ability of LLMs to transfer factual updates across languages. Even advanced models like Mistral and TowerInstruct, while effective in European languages, struggle significantly with low-resource languages. This limitation undermines the broader goal of making language technologies universally accessible and equitable Wang et al. (2023).
@@ Appendix A Model selection
Mistral-7B-Instruct-v0.2101010https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2: The model was developed by Jiang et al. (2023) and supports multilinguality [^11] https://encord.com/blog/mistral-large-explained/. It is designed around the causal language modeling framework. We shall refer to this model as Mistral.
TowerInstruct-7B-v0.2121212https://huggingface.co/Unbabel/TowerInstruct-7B-v0.2: This model Alves et al. (2024) has been developed on top of LLaMA2 Touvron et al. (2023) architecture and supports multilinguality including English, German, French, Spanish, Chinese, Portuguese, Italian, Russian, Korean, and Dutch. We shall refer to this model as TowerInstruct.
@@
Tamil-llama-7b-base-v0.1141414https://huggingface.co/abhinand/tamil-llama-7b-base-v0.1: This is a sophisticated model Balachandran (2023) developed specifically for bilingual tasks in Tamil and English, leveraging a 7 billion parameter causal language modeling framework. We shall refer to this model as Tamil-Llama.
@@ Appendix B Error analysis
Phonetic and Orthographic Errors These errors occur with words that sound similar (homophones) or are spelled similarly (homographs) but have different meanings. For instance, "their," "there," and "they’re" in English. Multilingual systems must accurately identify and apply the correct form based on context, a challenging task that often requires human-like understanding of language.

### v4

@@ 1 Introduction
Handling multilinguality in language models remains a significant challenge, particularly when models are prompted in languages other than English. Tasks such as question answering Xu et al. (2024a), addressing multilingual safety concerns Wang et al. (2024); Deng et al. (2024), or performing knowledge edits Hazra et al. (2024) often reveal noticeable gaps in performance for low-resource languages. Despite the advancements in multilingual large language models (LLMs), disparities persist, especially for languages with fewer computational resources. A clear example of this issue arises in knowledge editing Sinitsin et al. (2020); De Cao et al. (2021). For instance, when an LLM is updated to correct a factual statement, “The PM of the UK is Rishi Sunak" to “The PM of the UK is Keir Starmer" the model may apply the update accurately in well-represented languages like English or French Qi et al. (2023); Xu et al. (2023). However, the same edit often fails to propagate when queried in low-resourced languages like Tamil or Hindi. This inconsistency highlights a critical weakness in the ability of LLMs to transfer factual updates across languages. Even advanced models like Mistral and TowerInstruct, while effective in European languages, struggle significantly with low-resource languages. This limitation undermines the broader goal of making language technologies universally accessible and equitable Wang et al. (2023).
@@ 7 Error analysis
8 Discussion
Here we discuss two important questions – How do multilingual LLMs handle cross-lingual knowledge edits? and What steps can industry practitioners take to address cross-lingual disparities?
{mdframed}
[roundcorner=6pt,backgroundcolor=gray!10]
How do multilingual LLMs handle cross-lingual knowledge edits?
Modern LLMs often fail to propagate factual updates consistently across languages. While languages like English, French, and German benefit from extensive corpora Xu et al. (2024b), those like Hindi, Tamil, and Kannada suffer from data scarcity, causing unstable knowledge transfer Qi et al. (2023). Further, editing methods ROME and MEMIT encounter problems with highly agglutinative or morphologically rich languages.
Key observations
  • Data scarcity: Inadequate corpora produce sparse embeddings, disrupting the model’s ability to adapt newly introduced facts Das et al. (2022).
  • Architectural bias: LLM pipelines typically prioritize English, overlooking morphological idiosyncrasies in languages like Tamil or Kannada.
  • Complex linguistic features: Idiomatic expressions and cultural references can invalidate edits that were accurate in English Beniwal et al. (2024); merging specialized models can exacerbate divergences if representations are misaligned Yadav et al. (2023).
{mdframed}
[roundcorner=6pt,backgroundcolor=gray!5]
What steps can industry practitioners take to address cross-lingual disparities?
A holistic approach is needed to ensure consistent, multi-lingual fact-editing. Below are five key strategies:
  • Expand low-resource corpora:
    Rationale: Larger, more representative datasets address embedding sparsity;
    Implementation: Generate crowdsourced/synthetic data Hazra et al. (2024).
  • Continuous model editing:
    Rationale: Iterative edits balance new knowledge with existing facts [^10] https://www.microsoft.com/en-us/research/blog/lifelong-model-editing-in-large-language-models-balancing-low-cost-targeted-edits-and-catastrophic-forgetting/; primarily important for industries dealing with finance, healthcare, and law (e.g., updating a multilingual LLM to reflect new data privacy laws (GDPR, CPRA) in different regions without retraining from scratch).
    Case study: Microsoft’s lifelong editing merges local patches with broader retraining Cao et al. (2021).
  • Alignment-focused architectures:
    Rationale: Combine morphological analysis, advanced NER, & cross-lingual parameter sharing;
    Benefit: Stable knowledge propagation in structurally diverse languages Wang et al. (2023).
  • Dedicated edit modules:
    Rationale: Log each update & validate in all languages to avoid accidental overwrites;
    Implementation: Use an “edit ledger” in attention layers Hase et al. (2023).
  • Rigorous multilingual testing:
    Rationale:
    Systematic checks prevent bias & misinformation from creeping in;
    Tools: Curated test suites for reliability, cultural fitness, and domain-specific accuracy Hazra et al. (2024).
@@ Appendix A Model selection
Mistral-7B-Instruct-v0.2111111https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2: The model was developed by Jiang et al. (2023) and supports multilinguality [^12] https://encord.com/blog/mistral-large-explained/. It is designed around the causal language modeling framework. We shall refer to this model as Mistral.
TowerInstruct-7B-v0.2131313https://huggingface.co/Unbabel/TowerInstruct-7B-v0.2: This model Alves et al. (2024) has been developed on top of LLaMA2 Touvron et al. (2023) architecture and supports multilinguality including English, German, French, Spanish, Chinese, Portuguese, Italian, Russian, Korean, and Dutch. We shall refer to this model as TowerInstruct.
@@
Tamil-llama-7b-base-v0.1151515https://huggingface.co/abhinand/tamil-llama-7b-base-v0.1: This is a sophisticated model Balachandran (2023) developed specifically for bilingual tasks in Tamil and English, leveraging a 7 billion parameter causal language modeling framework. We shall refer to this model as Tamil-Llama.
@@ Appendix B Error analysis
Phonetic and orthographic errors These errors occur with words that sound similar (homophones) or are spelt similarly (homographs) but have different meanings. For instance, "their," "there," and "they’re" in English. Multilingual systems must accurately identify and apply the correct form based on context, a challenging task that often requires human-like understanding of language.